# Exploring Weak measurements within the Einstein-Dirac Cosmological framework

### v1

@@ II The Einstein-Dirac Solution in Friedmann-Lemaître-Robertson-Walker Space
$\left\{\begin{array}[]{lll}{w}^{1}&=&-\frac{R_{max}}{m}\\ {w}^{2}&=&\sqrt{N^{2}-\frac{R^{2}_{max}}{m^{2}}}\cos(2mt+\phi_{\text{max}}-2mt_{\text{max}})\\ {w}^{3}&=&\sqrt{N^{2}-\frac{R^{2}_{max}}{m^{2}}}\sin(2mt+\phi_{\text{max}}-2mt_{\text{max}}),\end{array}\right.$ (73)
where $m$ is the mass of Dirac particle.
Subtsituting $\lambda=0$ into the components of $\vec{w}$ in (31-33), we obtain
@@
$\left\{\begin{array}[]{lll}{w}^{1}&=&-\left(\rho_{out_{1}}^{2}-\rho_{out_{2}}^{2}\right)\\ {w}^{2}&=&2\rho_{out_{1}}\rho_{out_{2}}\sin(\eta_{out_{1}}-\eta_{out_{2}})\\ {w}^{3}&=&2\rho_{out_{1}}\rho_{out_{2}}\cos(\eta_{out_{1}}-\eta_{out_{2}}).\end{array}\right.$ (81)
@@
Figure 7: The acceleration of the scale function in an Universe dominated by matter.
@@ III Weak measurement on the Einstein-Dirac solutions
The general formula of a weak measurement of any observable is given by
@@
A pictorial description of a weak measurement in our context is like in figure 8.
@@
Because of the time dependency of the scale function, an exact evolutionary operator is impossible to compute. Still, it can be approximated by using the Wentzel–Kramers–Brillouin (WKB) method. Following Finster [29], the smallness of the Compton wavelength compared to the lifetime of the Universe justifies the use of WKB-type approximation. Therefore, we define the approximated unitary evolutionary operator as
@@
Let us point out here that $U(t,t_{\text{in}})$ is given by
$U^{-1}(t)\left(\begin{array}[]{cc}e^{-iF(t,t_{\text{in}})}&0\\ 0&e^{iF(t,t_{\text{in}})}\end{array}\right)U(t_{in})$
wherein the preselected state is taken from the radiation-dominated Universe with $m=0$, which yields
$g(t_{\text{in}})=\frac{1}{2}\arctan(\frac{\lambda}{m*R(t_{\text{in}})})=\frac{\pi}{4}+k\pi$ for $m=0$. For $k=0$, the matrix will be given by
$U(t_{\text{in}})=\frac{\sqrt{2}}{2}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$
Whereas the postselected state is taken from the dust-dominated Universe wherein $\lambda=0$. This will yield
$g(t)=\frac{1}{2}\arctan(\frac{\lambda}{m*R(t_{\text{in}})})=0+k\pi$ for $\lambda=0$, and for $k=0$ this results in $U(tout)$ is the identity matrix
$U(t_{out})=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right).$
@@
The unitary operator approximated, we can now compute the weak measurement of some observable starting from the Energy-momentum Tensor.
@@
In this section, we compute the weak measurement of the energy-momentum tensor given by
@@
$G^{r}=\frac{f(r)}{R(t)}\left(\cos\vartheta\gamma^{3}+\sin\vartheta\cos\varphi\gamma^{1}+\sin\vartheta\sin\varphi\gamma^{2}\right),$ (92)
$G^{\vartheta}=\frac{1}{rR(t)}\left(-\sin\vartheta\gamma^{3}+\cos\vartheta\cos\varphi\gamma^{1}+\cos\vartheta\sin\varphi\gamma^{2}\right),$ (93)
$G^{\varphi}=\frac{1}{rR(t)\sin\vartheta}\left(-\sin\varphi\gamma^{1}+\cos\varphi\gamma^{2}\right),$ (94)
@@
where $I$ is the $2\times 2$ identity matrix and $\sigma^{\alpha}$ are the Pauli matrices:
$\sigma^{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\sigma^{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\quad\sigma^{3}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
The differential operator is given by $D_{\mu}=\partial_{\mu}-iE_{\mu}\text{, }$ and $G^{\mu}$ obeys to the anti-commutation rule.
@@
To compute the numerator part of the weak measurement of the energy-momentum tensor, following Finster in[29], we adapt the object $P(t,\vec{x};t^{\prime},\vec{x^{\prime}})$ to our need
@@
wherein $E_{\lambda}$ denote the spectral projectors of $\mathcal{D}_{S^{3}}$, given by $E_{\pm\mid\lambda\mid}=\displaystyle\sum^{\mid\lambda\mid-\frac{3}{2}}_{n=0}\sum^{j}_{k=-j}\psi^{\pm}_{njk}(x)\overline{\psi^{\pm}_{njk}(\acute{x})}$. Then, the numerator part of the energy-momentum tensor, $\langle\psi_{\text{out}}|T_{\mu\nu}|\psi_{\text{in}}\rangle(t,\vec{x})$ can be expressed in terms of $P$ by
@@
where j represents $r,\vartheta,\varphi$. Due to spherical symmetry and homogeneity of the space, $T_{r}^{r}=T_{\vartheta}^{\vartheta}=T_{\varphi}^{\varphi}$ are all equal.
One sees here that if $|\psi_{in}(t)\rangle=|\psi_{out}(t)\rangle$, the weak measurement of the energy-momentum tensor will coincide with the one computed in [1]. In this regard, weak measurements are generalizations of classical ones.
@@
For better analysis, we construct a complex Bloch vector since it is constructed in the same manner as a Bloch vector, but the components are complex functions. Aharonov and Gruss constructed a density operator from the Two-State vectors in their paper Two-time interpretations of quantum mechanic [9]. This complex Bloch vector is linked to that density operator. One should note that if the preselected state is equal to the postselected one, then we shall get the same construction as in the paper of Finster and Hainzl on page 14 in [1].
@@
As it is difficult to have a simple expression of the weak measurement of the energy-momentum tensor to make an interpretation, let us compute a concrete example with values that make the computation easy. Let us consider $\lambda=3/2$, the eigenvalue of the Dirac operator $\mathcal{D}_{\mathcal{H}}$, $t_{\text{in}}\approxeq 0$ and $R_{0}=10^{-3}$. Theses conditions imply $G(t_{\text{in}})=0$ in (50). They will yield the following expression: $\rho_{in_{1}}\approx 2$ and $\rho_{in_{2}}\approx 0$ and $\eta_{in_{1}}-\eta_{in_{2}}=0$. For simplicity, we shall consider $\eta_{in_{1}}=\eta_{in_{2}}=0$.
We also remind that the preselected state is taken from the radiation-dominated Universe with $m=0$ which yields $g(t_{\text{in}})=\frac{1}{2}\arctan(\frac{\lambda}{0*R(t_{\text{in}})})=\frac{\pi}{4}+k\pi$. For $k=0$, the matrix will be given by $U(t_{\text{in}})=\frac{\sqrt{2}}{2}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$. Therefore the preselected state at $t_{\text{in}}$ will be $\alpha_{\text{in}}(t_{\text{in}})=2$ and $\beta_{\text{in}}(t_{\text{in}})=0$.
As for the postselected state vector, the scale function and the time expressed $R(g)=\frac{R_{\text{max}}}{2}\left(1-\cos(g)\right)$, $t(g)=\frac{R_{\text{max}}}{2}\left(g-\sin(g)\right)$, we have the following observations
  • A maximal $R(g)$ will be $R_{\text{max}}$, and the maximal $t$, that is $t_{\text{out}}$ will be $R_{\text{max}}k\frac{\pi}{2}$.
@@
We remind that the postselected state is taken from the dust-dominated Universe wherein $\lambda=0$. This will yield $g(t)=\frac{1}{2}\arctan(\frac{0}{m*R(t_{\text{in}})})=0+k\pi$ This will lead, for $k=0$, to $U(t_{out})=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$, an identity matrix.
The postselected state $t_{\text{out}}$ will be
@@
For simplicity, let us consider a weak measurement taking place at $t=t_{out}$ Universe and $R_{\text{max}}=10$ and $m=21.5$. This implies that
@@
$U(t_{\text{out}},t_{\text{in}})=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}e^{-i\pi m\text{Rmax}}&-e^{-i\pi m\text{Rmax}}\\ e^{i\pi m\text{Rmax}}&e^{i\pi m\text{Rmax}}\\ \end{array}\right),$
@@
$\begin{split}&\alpha_{\text{in}}(t_{\text{out}})=\frac{1}{\sqrt{2}}(\alpha_{\text{in}}-\beta_{\text{in}})e^{-i\pi m\text{Rmax}},\\ &\beta_{\text{in}}(t_{\text{out}})=\frac{1}{\sqrt{2}}(\alpha_{\text{in}}+\beta_{\text{in}})e^{i\pi m\text{Rmax}}.\end{split}$ (106)
Since the weak measurement is taking place at $t=t_{\text{out}}$, the postselected stated is fixed.
@@
$\begin{split}&v_{f,i}^{0}=1.6e^{i\pi mR_{\text{max}}}-(2.3i)e^{i\phi_{\text{max}}-i\pi mR_{\text{max}}}\\ &v_{f,i}^{1}=(2.3i)e^{i(\pi mR_{\text{max}}+\phi_{\text{max}})}+1.6e^{-i\pi mR_{\text{max}}}\\ &v_{f,i}^{3}=(2.3i)\left(-e^{i\phi_{\text{max}}-i\pi mR_{\text{max}}}\right)-1.6e^{i\pi mR_{\text{max}}}\end{split}$ (108)
@@
In (24) the acceleration of the Universe is given in terms of the energy momentum tensor. In (29), the acceleration of the Universe is given in terms of the Bloch vector. But in the Two-State Formalism, the Hopf transformation yields a complex Bloch vector. The Two-State Formalism generalizes the One-State Formalism since it suffices to consider the same state vector in the Two-State Formalism to have the classical Bloch vector. We propose to give here the acceleration of the universe in terms of weak measurements of the energy-momentum tensors and derive it in terms of Complex Bloch vector. The acceleration of the Universe would then be given as
@@
In this formula, one clearly sees that the acceleration of the Universe may be comprehended from the Two-State Formalism and weak measurements theories. Assuming a different postselected state vector may change the shape of the acceleration of the Universe. We remind here that the real and the imaginary parts of a weak value can be measured. The real part of a weak value may be interpreted as the conditioned average associated with an observable in Two Vector Formalism, [24], [32], and they represent the shift of the average detected position due to postselection. The imaginary part of the weak value represents the shift of the average impulsion due to postselection.
The computation of the acceleration of the Universe with the above condition would yield
@@
This example shows that the acceleration is not Zero and depends here on the postselection.
@@
The difference between Figure 9 and 10 shows how the acceleration of the Universe may change with a different postselection. We have postselected a state vector with $\phi_{\text{max}}=0$ in Figure 9 and $\phi_{\text{max}}=\frac{\pi}{4}$ in Figure 10. It is clear that, this difference causes more acceleration in Figure 10 than in Figure 9. In conclusion, we have shown how to compute the weak measurement of the Energy-momentum tensor. We derived how to amplify weak values of the energy-momentum tensor using an appropriate state vector almost orthogonal to the preselected one. We have finally derived the weak values of the acceleration of the Universe and have showed how sensitive it is compared to classical acceleration.
III.2 Weak measurement of $\sigma^{z}$ operator.
@@
Following Kofman, Cormann and Ferraz in [33], [34] and [7] respectively, the weak measurement of the generator of $SU(2)$ is given by
@@
In our case, $\vec{r}=(\begin{array}[]{ccc}0&0&1\end{array})$ to get $\sigma^{z}$. Following Kofman in [33], the vectors $\vec{f}$ and $\vec{i}$ are computed as follows
@@
In the same way, the preselected state is given by:
@@
For simplicity we shall consider $t_{\text{in}}=0$ which implies that $\eta_{\text{in1}}-\eta_{\text{in2}}=0$. We will set $\eta_{in_{1}}-\eta_{in_{2}}=\Delta_{in}$ when necessary. This implies $\Delta_{\text{in}}=0$ and $t_{\text{out}}=t_{\text{max}}$. We obtain the following Postselected and Preselected vectors $\vec{w_{f}}$ and $\vec{w_{i}}$ whose components are given as $\vec{w_{f}}=$
$\left\{\begin{array}[]{lll}&{w_{f}}^{1}=-\frac{R_{\text{max}}}{m}\sin{(2g(t))}+\\ &\sqrt{N^{2}-\frac{R^{2}_{max}}{m^{2}}}\cos{\left(\frac{\pi}{2}-\phi_{\text{max}}-2F(t_{\text{out}},t)\right)}\cos{(2g(t))}\\ &{w_{f}}^{2}=\sqrt{N^{2}-\frac{R^{2}_{max}}{m^{2}}}\sin{(\frac{\pi}{2}-\phi_{\text{max}}-(2F(t_{\text{out}},t)))}\\ &{w_{f}}^{3}=\frac{R_{\text{max}}}{m}\cos{(2g(t))}+\\ &\sqrt{N^{2}-\frac{R^{2}_{max}}{m^{2}}}\cos{(\frac{\pi}{2}-\phi_{\text{max}}-2F(t_{\text{out}},t))}\sin{(2g(t))}\\ \end{array}\right.,$ (122)
and $\vec{w_{i}}=$

### v2

@@ II The Einstein-Dirac Solution in Friedmann-Lemaître-Robertson-Walker Space
$\left\{\begin{array}[]{lll}{w}^{1}&=&-\frac{R_{max}}{m}\\ {w}^{2}&=&\sqrt{N^{2}-\frac{R^{2}_{max}}{m^{2}}}\cos{\left(\Phi_{out}(t)\right)}\\ {w}^{3}&=&\sqrt{N^{2}-\frac{R^{2}_{max}}{m^{2}}}\sin{\left(\Phi_{out}(t)\right)},\end{array}\right.$ (72)
where
$\Phi_{out}(t)=2mt+\phi_{\text{max}}-2mt_{\text{max}}$ (73)
$m$ is the mass of Dirac particle and $\phi_{\text{max}}$ a phase.
Substituting $\lambda=0$ into the components of $\vec{w}$ in (29-31), we obtain
@@
$\left\{\begin{array}[]{lll}{w}^{1}&=&-\left(\rho_{out_{1}}^{2}-\rho_{out_{2}}^{2}\right),\\ {w}^{2}&=&2\rho_{out_{1}}\rho_{out_{2}}\sin(\eta_{out_{1}}-\eta_{out_{2}}),\\ {w}^{3}&=&2\rho_{out_{1}}\rho_{out_{2}}\cos(\eta_{out_{1}}-\eta_{out_{2}}).\end{array}\right.$ (81)
@@ III Weak measurement on the Einstein-Dirac solutions
A weak value of any observable is given by
@@
Weak values are results of weak measurements. They are also derivable independently from weak measurements from the classical expectation values in classical quantum mechanics [33]. For instance, one can derive it from the expectation values of an observable and see it as intimately linked to the weak values. If $\hat{A}$ is an observable, then
$\langle\psi|\hat{A}|\psi\rangle$ $=\langle\psi|\hat{A}\int|\phi\rangle\langle\phi|d\phi|\psi\rangle$
[…]
wherein $\langle\hat{A}\rangle_{w}=\frac{\langle\psi|\hat{A}|\phi\rangle}{\langle\psi|\phi\rangle}$ is the so called weak value of the observable $\hat{A}$ and $P_{\phi\psi}=|\langle\phi|\psi\rangle|^{2}$ is the detection probability of event $\phi$ out of the prepared initial state $\psi$. This probability is that of measuring $\langle\hat{A}\rangle_{w}$.
An alternative way is to see $|\psi\rangle$ as a sum or difference of two other state vectors, let say $|\psi\rangle=|\psi_{1}\rangle\pm|\psi_{2}\rangle$. Then we will have
$\langle\psi|\hat{A}|\psi\rangle$ $=(\langle\psi_{1}|\pm\langle\psi_{2}|)\hat{A}(|\psi_{1}\rangle\pm|\psi_{2}\rangle)$
$=\langle\psi_{1}|\hat{A}|\psi_{1}\rangle+\langle\psi_{2}|\hat{A}|\psi_{2}\rangle\pm 2Re\langle\psi_{1}|\hat{A}|\psi_{2}\rangle$
$=\langle\psi_{1}|\hat{A}|\psi_{1}\rangle+\langle\psi_{2}|\hat{A}|\psi_{2}\rangle\pm 2Re\frac{\langle\psi_{1}|\hat{A}|\psi_{2}\rangle}{\langle\psi_{1}|\psi_{2}\rangle}\langle\psi_{2}|\psi_{1}\rangle$
$=\langle\hat{A}_{1}\rangle+\langle\hat{A}_{2}\rangle\pm 2Re\langle\hat{A}_{1,2}\rangle_{w}\langle\psi_{2}|\psi_{1}\rangle$
wherein $\langle\hat{A}_{1}\rangle=\langle\psi_{1}|\hat{A}|\psi_{1}\rangle$, $\langle\hat{A}_{2}\rangle=\langle\psi_{2}|\hat{A}|\psi_{2}\rangle$ and $\langle\hat{A}_{1,2}\rangle_{w}=\frac{\langle\psi_{1}|\hat{A}|\psi_{2}\rangle}{\langle\psi_{1}|\psi_{2}\rangle}$.
In (85), state vectors are multiplied by evolutionary operators. A pictorial description of weak measurement in our context is represented in figure 6.
@@
Because of the complexity involving the construction of the exact evolutionary operator due to the time dependency of the Dirac Hamiltonian, we are going to approximate it by using the Wentzel–Kramers–Brillouin (WKB) method. Also, following Finster [30], the smallness of the Compton wavelength compared to the lifetime of the Universe justifies the use of WKB-type approximation. The approximated unitary evolutionary operator will be expressed as
@@
The unitary operator applied to the pre-selected state $U(t,t_{\text{in}})$ is given as
$U^{-1}(t)\left(\begin{array}[]{cc}e^{-iF(t,t_{\text{in}})}&0\\ 0&e^{iF(t,t_{\text{in}})}\end{array}\right)U(t_{in}).$
Since the pre-selected state is taken from the radiation-dominated Universe where $m=0$, this will yield
$g(t_{\text{in}})=\frac{\pi}{4}+k\pi$. For $k=0$,
$g(t_{\text{in}})=\frac{\pi}{4}$ (89)
and the matrix will be given by
$U(t_{\text{in}})=\frac{\sqrt{2}}{2}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}.$ (90)
Whereas the post-selected state is taken from the dust-dominated Universe wherein $\lambda=0$. This will yield
$g(t_{out})=0+k\pi$ for $\lambda=0$, and for $k=0$
$g(t_{\text{in}})=0$ (91)
and the unitary matrix $U(t_{out})$ is the identity matrix
$U(t_{out})=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right).$ (92)
@@
The unitary operator approximated, we can now compute the weak value of some observables starting with the energy-momentum tensor.
@@
In this section, we compute the weak value of the energy-momentum tensor given by
@@
$G^{r}=\frac{f(r)}{R(t)}\left(\cos{\theta}\,\gamma^{3}+\sin{\theta}\cos{\phi}\,\gamma^{1}+\sin{\theta}\sin{\phi}\,\gamma^{2}\right),$ (96)
$G^{\theta}=\frac{1}{rR(t)}\left(-\sin{\theta}\,\gamma^{3}+\cos{\theta}\cos{\phi}\,\gamma^{1}+\cos{\theta}\sin{\phi}\,\gamma^{2}\right),$ (97)
$G^{\phi}=\frac{1}{rR(t)\sin{\theta}}\left(-\sin{\phi}\,\gamma^{1}+\cos{\phi}\,\gamma^{2}\right),$ (98)
@@
where $I$ is the $2\times 2$ identity matrix and $\sigma^{\alpha}$ are the Pauli matrices define in(7).
The differential operator is given by $D_{\mu}=\partial_{\mu}-iE_{\mu}\text{, }$ and $G^{\mu}$ obeys to the anti-commutation rule
@@
To compute the numerator part of the weak value of the energy-momentum tensor, following Finster in[30], we adapt the object $P(t,\vec{x};t^{\prime},\vec{x^{\prime}})$ to our need
@@
wherein $E_{\lambda}$ denote the spectral projectors of $\mathcal{D}_{S^{3}}$, given by $E_{\pm\mid\lambda\mid}=\displaystyle\sum^{\mid\lambda\mid-\frac{3}{2}}_{n=0}\sum^{j}_{k=-j}\psi^{\pm}_{njk}(x)\overline{\psi^{\pm}_{njk}(\acute{x})}$. Then, the numerator part of weak value of the energy-momentum tensor, $\langle\psi_{\text{out}}|T_{\mu\nu}|\psi_{\text{in}}\rangle(t,\vec{x})$ can be expressed in terms of $P$ by
@@
where j represents $r,\theta,\phi$. The space diagonal energy momentum tensor are all equal, that is $T_{r}^{r}=T_{\theta}^{\theta}=T_{\phi}^{\phi}$.
One sees here that if $|\psi_{in}(t)\rangle=|\psi_{out}(t)\rangle$, the weak value of the energy-momentum tensor will coincide with the one computed in [1]. In this regard, weak measurements are generalizations of classical averages.
@@
For better analysis, we construct a complex Bloch vector. Aharonov and Gruss constructed a density operator from the Two-State vectors in their paper Two-time interpretations of quantum mechanic [4]. This complex Bloch vector is linked to that density operator. One should note that if the pre-selected state is equal to the post-selected one, then we shall get the same construction as in the paper of Finster and Hainzl on page 14 in [1]. Our complex Bloch vector will be given as
@@
As it is difficult to have a simple expression of the weak value of the energy-momentum tensor to make an interpretation, let us calculate a concrete example using values that simplify the computation. For the pre-selected states let us have
[…]
These values will yield
[…]
Therefore the pre-selected state at $t_{\text{in}}$ will be given by
$\begin{cases}\alpha_{\text{in}}(t_{\text{in}})=\frac{\sqrt{2}}{2}\sqrt{2+2\cos\left(\Phi_{in}(t_{\text{in}})\right)}e^{-i\arctan{\left(\frac{3\sin{\left(\Phi_{in}(t_{\text{in}})\right)}}{R^{2}_{0}}\right)}}\\ \beta_{\text{in}}(t_{\text{in}})=\frac{\sqrt{2}}{2}\sqrt{2-2\cos\left(\Phi_{in}(t_{\text{in}})\right)}\end{cases}$ (111)
As for the post-selected state vector, the scale function and the time expressed as $R(\tau)=\frac{R_{\text{max}}}{2}\left(1-\cos(\tau)\right)$, $t(\tau)=\frac{R_{\text{max}}}{2}\left(\tau-\sin(\tau)\right)$ respectively, we have the following observations
  • A maximal $R(\tau)$ will be $R_{\text{max}}$, and the maximal $t$, that is $t_{\text{out}}$ will be $R_{\text{max}}k\frac{\pi}{2}$.
@@
We remind that the post-selected state is taken from the dust-dominated Universe wherein $\lambda=0$. This will yield $g(t_{out})=\frac{1}{2}\arctan(\frac{\lambda}{m*R(t_{\text{out}})})=0+k\pi$ This will lead, for $k=0$, to $U(t_{out})=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$, an identity matrix.
The post-selected state $t_{\text{out}}$ will be
@@
For simplicity, let us consider a weak measurement taking place at time $t_{out}$ and $R_{\text{max}}=10$ and $m=21.5$. This implies that
@@
$U(t_{\text{out}},t_{\text{in}})=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}e^{-i\pi mR_{\text{max}}}&-e^{-i\pi mR_{\text{max}}}\\ e^{i\pi mR_{\text{max}}}&e^{i\pi mR_{\text{max}}}\\ \end{array}\right),$
@@
$\begin{split}&\alpha_{\text{in}}(t_{\text{out}})=\frac{\sqrt{2}}{2}(\alpha_{\text{in}}(t_{\text{in}})-\beta_{\text{in}}(t_{\text{in}}))e^{-i\pi mR_{\text{max}}},\\ &\beta_{\text{in}}(t_{\text{out}})=\frac{\sqrt{2}}{2}(\alpha_{\text{in}}(t_{\text{in}})+\beta_{\text{in}}(t_{\text{in}}))e^{i\pi mR_{\text{max}}}.\end{split}$ (113)
Since the weak measurement is taking place at time $t_{\text{out}}$, the post-selected state is fixed.
@@
$\begin{split}&v_{f,i}^{0}=\text{C}\left((A-B)\cos{(\frac{\Phi_{in}}{2})}\text{D}+(A+B)\sin{(\frac{\Phi_{in}}{2})}\right)\\ &v_{f,i}^{1}=\text{C}(AB+1)\left(\cos{(\frac{\Phi_{in}}{2})}D-\sin{(\frac{\Phi_{in}}{2})}\right)\\ &v_{f,i}^{3}=C\left(\left(A+B\right)\cos{(\frac{\Phi_{in}}{2})}D+\left(A-B\right)\sin{(\frac{\Phi_{in}}{2})}\right)\end{split}$ (115)
wherein we have set $A=e^{2i\pi m\text{Rmax}}$ , $B=e^{i\text{$\phi$max}}$ and $C=e^{-i\pi m\text{Rmax}}\sqrt{\frac{\text{Rmax}}{m}+2}$, $D=e^{i\tan^{-1}\left(\frac{3\sin(\text{$\phi$tin})}{\text{R0}^{2}}\right)}$.
If we set $\Phi_{in}$ such that $\Phi_{in}=\pi$, we will have
$\begin{split}&v_{f,i}^{0}=\text{C}(A+B)\\ &v_{f,i}^{1}=-\text{C}(AB+1)\\ &v_{f,i}^{3}=C\left(A-B\right)\end{split}$ (116)
Such configuration happens if $t_{\text{in}}=t_{B}$ and $\phi_{0}=\pi$.
@@
In (22), the acceleration of the Universe is given in terms of the energy momentum tensor. In (27), it is given in terms of the Bloch vector. But in the Two-State Formalism, the Hopf transformation yields a complex Bloch vector. The Two-State Formalism generalizes the One-State Formalism since it suffices to consider the same state vector in the Two-State Formalism to have the classical Bloch vector. We propose to express the acceleration of the universe in terms of weak values of the energy-momentum tensors and derive it in terms of Complex Bloch vector. The acceleration of the Universe will then be given as
@@
In this formula, one clearly sees that the acceleration of the Universe may be comprehended from the Two-State Formalism and weak measurements theories. Assuming a different post-selected state vector may change the shape of the acceleration of the Universe. We remind here that the real and the imaginary parts of a weak value can be measured. The real part of a weak value may be interpreted as the best estimation of the conditioned average associated with an observable in Two Vector Formalism, [24], [34], and they represent the shift of the average detected position due to post-selection. The imaginary part of the weak value represents the shift of the average impulsion due to post-selection.
The computation of the acceleration of the Universe with the above condition yields
@@
This example shows that the acceleration is not Zero and depends here on the post-selection.
@@
The difference between Figure 7 and 8 shows how the acceleration of the Universe may change with a different post-selection. We have post-selected a state vector with $\phi_{\text{max}}=0$ in Figure 7 and $\phi_{\text{max}}=\frac{\pi}{4}$ in Figure 8. It is clear that, this difference causes more acceleration in Figure 8 than in Figure 7. In conclusion, we have shown how to compute the weak value of the Energy-momentum tensor. We derived how to amplify weak values of the energy-momentum tensor using an appropriate state vector almost orthogonal to the pre-selected one. We have finally derived the weak values of the acceleration of the Universe and have showed how sensitive it is compared to classical acceleration.
III.2 Weak value of $\sigma^{z}$ operator.
@@
Following Ferraz, Kofman and Cormann in [8, 35, 36] respectively, the weak value of the generator of $SU(2)$ is given by
@@
In our case, $\vec{r}=(\begin{array}[]{ccc}0&0&1\end{array})$. Following Kofman in [35], the vectors $\vec{f}$ and $\vec{i}$ are computed as follows
@@
We remind that $\alpha_{\text{out}}(t_{\text{out}})=\rho_{\text{out}_{1}}e^{i\eta_{\text{out}_{1}}}$ and $\beta_{\text{out}}(t_{\text{out}})=\rho_{\text{out}_{2}}e^{i\eta_{\text{out}_{2}}}$ wherein
$\rho_{\text{out}_{1}}$ $=\frac{\sqrt{2}}{2}\sqrt{N+\frac{R_{max}}{m}},$
$\rho_{\text{out}_{2}}$ $=\frac{\sqrt{2}}{2}\sqrt{N-\frac{R_{max}}{m}},$
$\eta_{\text{out}_{1}}-\eta_{\text{out}_{2}}$ $=\frac{\pi}{2}-\Phi_{out}(t_{\text{out}}),$
and
$\Phi_{out}(t_{\text{out}})=2mt_{\text{out}}+\phi_{\text{max}}-2mt_{\text{max}}$
In the same way, the pre-selected state is given by:
@@
Let us remind that $\alpha_{\text{in}}(t_{\text{in}})=\rho_{\text{in}_{1}}e^{i\eta_{\text{in}_{1}}}$ and $\beta_{\text{in}}(t_{\text{in}})=\rho_{\text{in}_{2}}e^{i\eta_{\text{in}_{2}}}$ wherein
$\rho_{\text{in}_{1}}$ $=\frac{\sqrt{2}}{2}\sqrt{N+\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos{\left(\Phi_{in}(t_{\text{in}})\right)}},$
$\rho_{\text{in}_{2}}$ $=\frac{\sqrt{2}}{2}\sqrt{N-\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos{\left(\Phi_{in}(t_{\text{in}})\right)}},$
$\eta_{in_{1}}-\eta_{in_{2}}$ $=-\arctan\left(\frac{\sqrt{\lambda^{2}N^{2}-R^{4}_{0}}\sin{\left(\Phi_{in}(t_{\text{in}})\right)}}{R^{2}_{0}}\right),$
and
$\Phi_{in}(t_{\text{in}})=4\lambda G(t_{\text{in}})-\lambda\pi-\phi_{0},$
and $\text{and }G(t_{\text{in}})=\tan^{-1}\left(\frac{t_{\text{in}}}{R(t_{\text{in}})}\right).$
We obtain the following post-selected $\vec{w_{f}}$ whose components are given as $\vec{w_{f}}=$
$\left\{\begin{array}[]{lll}&{w_{f}}^{1}=-\frac{R_{\text{max}}}{m}\sin{(2g(t))}+A\cos{\left(\Delta_{\text{out}}-2F(t_{\text{out}},t)\right)}\cos{(2g(t))},\\ &{w_{f}}^{2}=A\sin{(\Delta_{\text{out}}-2F(t_{\text{out}},t))},\\ &{w_{f}}^{3}=\frac{R_{\text{max}}}{m}\cos{(2g(t))}+A\cos{(\Delta_{\text{out}}-2F(t_{\text{out}},t))}\sin{(2g(t))},\\ \end{array}\right.,$ (128)
wherein we have set for readability $A=\left(N^{2}-\frac{R^{2}_{max}}{m^{2}}\right)^{\frac{1}{2}}$ and $\Delta_{\text{out}}=\eta_{\text{out}_{1}}-\eta_{\text{out}_{2}}$.
We obtain the pre-selected vector $\vec{w_{i}}$ whose components are given as $\vec{w_{i}}=$
@@
wherein we have considered the case $\Phi_{in}(t_{\text{in}})=0$. This is possible for $t_{in}=t_{B}$ (see (42) ) and $\phi_{0}=0$.